ConnersKitten:

Type the correct answer in each box. Use numerals instead of words. If necessary, use / for the fraction bar. Brenda is rolling a six-sided die and flipping a coin. Complete each of the following sentences regarding the probabilities involved in these actions using reduced fractions. The probability of rolling a five is . The probability of flipping the coin and it landing on heads is . The probability of rolling a five and landing on heads is . Are the events independent or dependent?

qwevery:

Answer: Experimental probability: 60% Which is 10% more than theoretical prob8 Step-by-step explanation: No. of odd outcomes: 12/20 = 0.6 0.6×100 = 60% Theoretical probability: 4/8 × 100 = 50% 60 - 50 = 10% more

ConnersKitten:

Thank you

qwevery:

np:)